OpenStudy (anonymous):

The change in water vapor in a cloud is modeled by a polynomial function, C(x). Describe how to find the x-intercepts of C(x) and how to construct a rough graph of C(x) so that the meteorologist can predict when there will be no change in the water vapor. You may create a sample polynomial to be used in your explanations.

OpenStudy (anonymous):

@phi can you help please

OpenStudy (kohai):

Ah, alright. If you have a function, you find the x-intercepts by factoring the function and then solving for x. For example, \[c(x)=x^2-4\] \[c(x)=(x+2)(x-2)\] \[x=2, -2\]

OpenStudy (anonymous):

ok

OpenStudy (kohai):

To graph it, all you have to do is plug some x-values into the equation. \[c(x)=x^2-4\] \[c(1)=1^2-4\] \[c(x)=x^2-4\] \[c(2)=2^2-4\] And you can plug a few more in and solve, and then graph the values you get.

OpenStudy (kohai):

Does that make sense?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

Did you use an example that couldnt be used for this problem?

OpenStudy (anonymous):

because I got a little lost during the first part

OpenStudy (kohai):

No, no it could be used. It was just an example. Here's another one, if you wish. \[c(x)=x^2+5x+6\] \[c(x)=(x+2)(x+3)\]

OpenStudy (anonymous):

Oh alright! I didnt see the part where you said to plug it in haha thanks!

OpenStudy (kohai):

You're very welcome :)